INFRARED DEPLETION SPECTROSCOPY OF THE ANILINE-WATER-TOLUENE CLUSTER CATION AND MODE SELECTIVE PREDISSOCIATION OF THE HYDROGEN BONDS.

NAVEED K. PIRACHA, \underline{TAISUKE NAKANAGA}, Photoreaction Control Research Center, National Institute of Advanced Industrial Science and Technology (AIST), Tsukuba, Ibaraki 305-8565, Japan.

The vibrational spectrum of the ternary cluster cation of aniline-water-toluene in a supersonic jet has been measured by infrared depletion method in the NH and OH stretching region. Three bands were observed at 3722, 3634 and 3295cm^{−1}. The analysis of the vibrational spectrum showed that two NH bonds of aniline interact with water and toluene in the cluster cation. The band at 3295cm^{−1} was assigned to the stretching vibration of NH of aniline cation interacting with toluene, whereas other two bands were assigned to the symmetric and anti-symmetric stretching vibrations of free OH of water. When the cluster absorbs the infrared photon, it dissociated into two fragments. There are two possible dissociation path of this cluster.

1) AWT^{+} → AW^{+} + T and 2) AWT^{+} → AT^{+} + W The branching ratio of this reaction AW^{+}/AT^{+} was measured for three different bands, and we obtained 0.49 for the NH stretching vibration and 0.37 and 0.38 for the OH stretching vibrations. This results suggest that there is a fast energy transfer from the NH stretching vibration to the hydrogen bond between NH and toluene which enables the vibrational mode dependence of the predissociation.